myininaya (myininaya):

Maybe I should rewrite my post: can someone pretty please write the first couple of terms for summation(2,i=-1..4)

myininaya (myininaya):

\[\sum_{i=-1}^{4}2\]

myininaya (myininaya):

i believe in you cyter :)

OpenStudy (anonymous):

2+2+2+2+2+2=12

myininaya (myininaya):

why is it that though?

OpenStudy (anonymous):

-1,0,1,2,3,4, are 6 items

myininaya (myininaya):

i might be losing my brain lol or something but why does that tell us to do 2(6)

OpenStudy (anonymous):

i think of it this way sum of 2(i)^0 from i=-1 to 4 = 2(-1)^0+ 2(0)^0+2(1)^0+2(2)^0+2(3)^0+2(4)^0 = 12 if you know something about arithmetic series it should help

myininaya (myininaya):

i know but that 0^0 is confusing me because it is undefined

OpenStudy (anonymous):

0^0 is indeterminate this is tricky here

myininaya (myininaya):

i have always thought of it just like you said but I never thought about the 0^0

OpenStudy (anonymous):

okay, scratch my method, and check this out http://betterexplained.com/articles/techniques-for-adding-the-numbers-1-to-100/

myininaya (myininaya):

i don't know if this helps

myininaya (myininaya):

I do thank you though for trying to help me i will give you a medal

OpenStudy (anonymous):

does that help? We are taking the sum of the y values for f[n]=n. at f[-1]+f[0]+...+f[4] a way of visualizing functions might help

myininaya (myininaya):

but we have f(n)=2 omg omg ok so we f(-1)+f(0)+f(1)+... =2+2+2+ omg you are my hero

myininaya (myininaya):

i never thought of visualing it as a function

OpenStudy (anonymous):

lol, good. Was just about to suggest the f(n) = 2 thing =) Glad you hit upon it.

myininaya (myininaya):

:)

myininaya (myininaya):

cyter eveytime i notice you post something i'm gonna give you a medal

myininaya (myininaya):

i just went back through some old posts and probably gave you like 6 medals lol

OpenStudy (anonymous):

Thanks :)

myininaya (myininaya):

like seriously cyter you are my hero

OpenStudy (anonymous):

Lets hope I save more... onto infinity!

OpenStudy (anonymous):

11 medals are enough

myininaya (myininaya):

i know alot of math. I didn't think anyone could really impress me on this site or save me.

OpenStudy (anonymous):

one was probably enough

OpenStudy (anonymous):

but its ok

myininaya (myininaya):

i mean i exaggerate when i say save lol

OpenStudy (anonymous):

I did exaggerate too

myininaya (myininaya):

i have a masters in math but there are small things i just accepted rules on

myininaya (myininaya):

its nice to understand

OpenStudy (anonymous):

This is the best thing I like about collaborative learning You get to learn and really reach levels of understanding you couldn't reach

myininaya (myininaya):

Polpak is pretty smart and impressive but no one has ever rescued me like you lol

myininaya (myininaya):

i'm sorry but i'm going to medal stalk you

OpenStudy (anonymous):

I appreciate it but I'm really interested in understanding concepts in different ways/dimensions thanks to people kind enough to enlighten everyone.

OpenStudy (anonymous):

I've also learned quite a bit from your answers

myininaya (myininaya):

lol well it is the only i can do probably to show thanks

OpenStudy (anonymous):

Your word is enough.

myininaya (myininaya):

have your really learned from my answers

myininaya (myininaya):

questions you asked?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm not sure

myininaya (myininaya):

what level of math are you?

OpenStudy (anonymous):

Calc 1

OpenStudy (anonymous):

my power had gone out so sorry about the delay

myininaya (myininaya):

omg seriously wow cal 1 omg you are a genius already

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

what math class r u?

myininaya (myininaya):

i have my masters in math lol

OpenStudy (anonymous):

u r the genius

myininaya (myininaya):

no! i couldn't explain that summation thing up there lol

myininaya (myininaya):

i mean now i can

myininaya (myininaya):

gn